# Birational Geometry of Special Quotient Foliations and Chazy’s Equations

### v1

@@ 7.2. The birational automorphisms of Brunella’s very special foliation
We will consider the quotient model for $\mathcal{F}_{3}$ described in Section 3.1. The foliation $\mathcal{E}_{3}$ given by (15) is the foliation $\mathcal{L}_{\rho}$ of Section 7.1 for the primitive sixth root of unity $\rho=-\omega^{2}$. The transformation $T_{3}$ of $\mathbb{P}^{2}$ in (16) is the monomial transformation (34) associated to $\left(\begin{array}[]{rr}0&-1\\ 1&-1\end{array}\right)$; its action on $\mathbb{H}$ via (33) fixes $\rho$.
@@
The action of $T_{3}$ on $\mathbb{P}^{2}$ preserves both $U$ and the foliation $\mathcal{L}_{\rho}$, and acts thus upon the corresponding leaf space. From (7.1), this last action is given by the order-three biholomorphism $T^{\flat}_{3}$ of $E_{\rho}$ induced by $z\mapsto-\rho z$, for which $f_{\tau}\circ T_{3}=T^{\flat}_{3}\circ f_{\tau}$. The action of $T_{3}$ on $\mathbb{P}^{2}$ multiplies $X_{\rho}$ by the constant $\rho^{2}$, and, in consequence, preserves the affine structure along the leaves of $\mathcal{L}_{\rho}$. This endows the leaves of $\mathcal{F}_{3}$ on the regular part of $U/T_{3}$ with an affine structure along each leaf, varying holomorphically (a foliated affine structure). The leaves of $\mathcal{L}_{\rho}$ that correspond to points in $E_{\rho}$ with trivial stabilizer under the action of $T^{\flat}_{3}$ map injectively to $\mathbb{P}^{2}/T_{3}$ as leaves of $\mathcal{F}_{3}$; their images are transcendental, have saturated neighborhoods that are injective images of covering tubes, and are without holonomy. A leaf of $\mathcal{L}_{\rho}$ that corresponds to a point in $E_{\rho}$ with non-trivial stabilizer under the action of $T_{3}$ (that is fixed by it) maps in a three-to-one ramified way to a leaf of $\mathcal{F}_{3}$ having a holonomy of order three (there are three such leaves); the affine structure in each one of these leaves is inherited from the one on $\mathbb{C}$ under the quotient by multiplication by $\rho^{2}$, with the point $0\in\mathbb{C}$ corresponding to a singular point of $\mathbb{P}^{2}/T_{3}$.
It will be convenient to consider the quotient $E_{\rho}/T^{\flat}_{3}$ as an orbifold, modeled on $\mathbb{P}^{1}$, with three conic points of angle $32\pi/$ 3. The space of leaves of $\mathcal{F}_{3}$ may be described as the one given by identifying transversals to the foliation under the holonomy relation. The three leaves with holonomy correspond to the conical points of the orbifold structure for $E_{\rho}/T^{\flat}_{3}$.
@@
Let $R:\mathbb{P}^{2}/T_{3}\dashrightarrow\mathbb{P}^{2}/T_{3}$ be a birational automorphism of $\mathcal{F}_{3}$. It permutes holomorphically the Zariski-dense leaves of $\mathcal{F}_{3}$, in a holonomy-preserving way. In particular, $R$ induces a biholomorphism of $E_{\rho}/T^{\flat}_{3}$ compatible with its orbifold structure. Through its action on the three conical points, the orbifold group of biholomorphisms of $E_{\rho}/T^{\flat}_{3}$ identifies to the group $S_{3}$ of permutations on three symbols. This group may be generated by the maps induced by the transformations $Q$ and $S$ of Eqs. (17) and (18) on the space of transcendental leaves of $\mathcal{F}_{3}$. In order to prove the first item of Theorem 5, we will establish that $R$ belongs to the group generated by the maps induced by $Q$ and $S$. Up to composing with an element of this group, we may suppose that $R$ acts trivially on the leaf space of $\mathcal{F}_{3}$. Our aim to establish that $R$ is the identity.
Let $\Pi_{3}:\mathbb{P}^{2}\to\mathbb{P}^{2}/T_{3}$ denote the quotient map. Let $L\subset U$ be a leaf of $\mathcal{L}_{\rho}$ that has a non-trivial stabilizer under the action of $T_{3}$. Let $p\in L$ be a point that is not fixed by $T_{3}$, and that is such that $R$ is a biholomorphism on a neighborhood of $\Pi_{3}(p)$. Consider a sufficiently small transversal $j:D\to U$ as before, with $j(0)=p$. Consider also the associated covering tube $\psi:D\times\mathbb{C}\to U$, and suppose that it is preserved by $T_{3}$. Let $D^{*}=D\setminus\{0\}$.
We claim that there exists a holomorphic map $\widetilde{R}_{0}^{*}:\psi(D^{*}\times\mathbb{C})\to\psi(D^{*}\times\mathbb{C})$ such that, in restriction to $D^{*}\times\mathbb{C}$,
(41) $\Pi_{3}\circ\widetilde{R}_{0}^{*}=R\circ\Pi_{3}.$
The fundamental group of $\psi(D^{*}\times\mathbb{C})$ is cyclic, and identifies to that of $D^{*}$ via the projection $D^{*}\times\mathbb{C}\to D^{*}$. The restriction of $\Pi_{3}$ to $\psi(D^{*}\times\mathbb{C})$ is a Galois covering map onto its image, with group of deck transformations generated by $T_{3}$. As in Section 7.1, the covering tubes around the holonomy-free leaves of $\mathcal{F}_{3}$ are mapped by $R$ to similar tubes, affinely in restriction to each leaf, as in formula (37). In particular, $R$ is a biholomorphism in restriction to the open subset of $U/T_{3}$ formed by the holonomy-free leaves. The map $R\circ\Pi_{3}:\psi(D^{*}\times\mathbb{C})\to\Pi_{3}(\psi(D^{*}\times\mathbb{C}))$ is thus holomorphic, and the sought map $\widetilde{R}_{0}^{*}$ from Eq. (41) is a lift of the latter. Its existence follows from the well-known criterion for the existence of lifts to covering maps (see [Hat02, Prop. 1.33]) and from the facts that (i) the action of $R$ on the leaf space of $\Pi_{3}(\psi(D^{*}\times\mathbb{C}))$ is trivial, and (ii) this leaf space carries all the information of its fundamental group. This establishes the claim. Observe that there are three possible lifts $\widetilde{R}_{0}$, differing by post-composition with powers of $T_{3}$.
Consider a sufficiently small ball $B\subset U$ around $p$. In restriction to $B$, $R\circ\Pi_{3}$ is a biholomorphism onto its image, and there is thus a lift $\widetilde{r}:B\to U$ such that, in restriction to $B$, $\widetilde{r}\circ\Pi_{3}=\Pi_{3}\circ R$. Up to the action of $T_{3}$, we may suppose that $\widetilde{r}$ agrees with $\widetilde{R}_{0}^{*}$ in the intersection of the domains where each of them is defined. We have that $\widetilde{R}_{0}^{*}$ is affine along the leaves of $\mathcal{L}_{\rho}$ within $\psi(D^{*}\times\mathbb{C})$. This is a closed condition, and thus $\widetilde{r}$ is affine along $L$ as well: the map defined by $\widetilde{R}_{0}^{*}$ and $\widetilde{r}$ on $\psi(D^{*}\times\mathbb{C})\cup B$ is affine in restriction to each leaf, like in formula (37). This implies that $\widetilde{R}_{0}^{*}$ extends to a biholomorphic map $\widetilde{R}_{0}:\psi(D\times\mathbb{C})\to\psi(D\times\mathbb{C})$ such that $R\circ\Pi_{3}=\Pi_{3}\circ\widetilde{R}_{0}$. In particular, we have established that $R$ is a biholomorphism in restriction to $U/T_{3}$.
We affirm that there exists a global lift of $R|_{U}\circ\Pi_{3}$, a map $\widetilde{R}:U\to U$, preserving $\mathcal{L}_{\rho}$, such that $R\circ\Pi_{3}=\Pi_{3}\circ\widetilde{R}$. Let $U^{*}\subset U$ be the set formed by the leaves that are not fixed by $T_{3}$. Since $R$ acts by mapping each leaf onto itself, and since the leaf space carries all the information of the fundamental group, there exists a lift as above in restriction to $U^{*}$. By grafting maps like $\widetilde{R}_{0}$ on neighborhoods of the leaves with holonomy, we can extend this lift to all of $U$.
@@
By Proposition 19 and its proof, $\widetilde{R}$ belongs to the flow of $X_{\rho}$. In order for such automorphism of $\mathbb{P}^{2}$ to be a lift from one of $\mathbb{P}^{2}/T_{3}$, it must normalize the group generated by $T_{3}$, this is, either $\widetilde{R}T_{3}\widetilde{R}^{-1}=T_{3}$, or $\widetilde{R}T_{3}\widetilde{R}^{-1}=T_{3}^{2}$. However, the last possibility may be discarded, for the actions of both sides of the equality on the singularities of $\mathcal{L}_{\rho}$ do not agree. Thus, $\widetilde{R}$ and $T_{3}$ commute. If $\widetilde{R}$ is given by the flow of $X_{\tau}$ in time $b$, $\widetilde{R}(x,y)=(x\mathrm{e}^{b\rho},y\mathrm{e}^{b})$, and the above commutativity condition reads
$\left(\frac{1}{y}\mathrm{e}^{-b},\frac{x}{y}\mathrm{e}^{b(\rho-1)}\right)=\left(\frac{1}{y}\mathrm{e}^{-b\rho},\frac{x}{y}\mathrm{e}^{b}\right).$
We conclude that $b=0$, that $\widetilde{R}$ is the identity, and that $R$ is the identity as well, establishing the result.
@@ 7.3. The birational automorphisms of $\mathcal{F}_{4}$
Consider the quotient model for $\mathcal{F}_{4}$ of Section 3.2. The foliation $\mathcal{L}_{\mathrm{i}}$ of Section 7.1 gives a birational model for the foliation $\mathcal{E}_{4}$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ given by (19): on $\mathbb{P}^{2}$, blow up the points $(1:0:0)$ and $(0:1:0)$, and then blow down the strict transform of the line originally joining them. The automorphism $T_{4}$ in (20) corresponds to the monomial transformation associated to $\left(\begin{array}[]{rr}0&1\\ -1&0\end{array}\right)$, that fixes $\mathrm{i}$ through its action on $\mathbb{H}$. This corresponds to the order-four automorphism $T^{\flat}_{4}$ of $E_{\mathrm{i}}$ induced by multiplication by $\mathrm{i}$. The quotient is an orbifold modeled on $\mathbb{P}^{1}$, with three conical points: one with angle $\pi$, and two with angle $\pi/2$. Its orbifold group of biholomorphisms is generated by an involution that fixes the first point and permutes those of the latter pair. The map $J$ of Eq. (21) gives a birational map of $\mathcal{L}_{\mathrm{i}}/J$ that, through its action on the leaf space, induces this involution.
The arguments are essentially those of the previous case. Let $R:(\mathbb{P}^{1}\times\mathbb{P}^{1})/T_{4}\dashrightarrow(\mathbb{P}^{1}\times\mathbb{P}^{1})/T_{4}$ be a birational automorphism of $\mathcal{L}_{\mathrm{i}}$. Up to composing with $J$, we may suppose that $R$ acts trivially on the leaf space of $\mathcal{F}_{4}$. As before, it may be lifted to a birational automorphism $\widetilde{R}$ of $\mathbb{P}^{1}\times\mathbb{P}^{1}$ that is actually holomorphic, that belongs to the flow of $X_{\mathrm{i}}$ and that commutes with $T_{4}$. A straightforward calculation shows that $\widetilde{R}$, and thus $R$, must be the identity. This establishes the second item of Theorem 5.
@@ 7.4. On the absence of birational automorphisms of $\mathcal{F}_{6}$
Consider the quotient model for $\mathcal{F}_{6}$ described in Section 3.3, and keep the objects introduced in Section 7.2. The foliation $\mathcal{E}_{6}$ is birationally equivalent to $\mathcal{L}_{\rho}$, and its birational automorphism $T_{6}$, to the monomial transformation induced by $\left(\begin{array}[]{rr}0&1\\ -1&1\end{array}\right)$. The latter fixes $\rho$ through its action on $\mathbb{H}$. This corresponds to the biholomorphism of $E_{\rho}$ induced by multiplication by $\rho$. The quotient is an orbifold modeled on $\mathbb{P}^{1}$, with three conical points, of angles $1/2$, $1/3$ and $1/6$ times $2\pi$. It has a trivial group of biholomorphisms.
Let $R:\mathbb{P}^{2}/T_{6}\dashrightarrow\mathbb{P}^{2}/T_{6}$ be a birational automorphism of $\mathcal{L}_{\rho}$. It acts trivially on the leaf space of $\mathcal{F}_{6}$, and, as before, it may be lifted to a birational automorphism $\widetilde{R}$ of $\mathbb{P}^{2}$ that belongs to the flow of $X_{\rho}$ and that commutes with $T_{6}$, that can be shown to be the identity. This establishes the third item of Theorem 5.

### v2

@@ 7.2. The birational automorphisms of Brunella’s very special foliation
We will now prove the part of Theorem 5 concerning $\mathcal{F}_{3}$: that its only birational symmetries are those exhibited in Section 3.1. We will consider the quotient model for $\mathcal{F}_{3}$ described that section. The foliation $\mathcal{E}_{3}$ given by (15) is the foliation $\mathcal{L}_{\rho}$ of Section 7.1 for the primitive sixth root of unity $\rho=-\omega^{2}$. The transformation $T_{3}$ of $\mathbb{P}^{2}$ in (16) is the monomial transformation (36) associated to $\left(\begin{array}[]{rr}0&-1\\ 1&-1\end{array}\right)$; its action on $\mathbb{H}$ via (35) fixes $\rho$.
@@
The action of $T_{3}$ on $\mathbb{P}^{2}$ preserves both $U$ and the foliation $\mathcal{L}_{\rho}$, and induces an action of $T_{3}$ on its leaf space. From (7.1), this last action is given by the order-three biholomorphism $T^{\flat}_{3}$ of $E_{\rho}$ induced by $z\mapsto-\rho z$, for which $f_{\tau}\circ T_{3}=T^{\flat}_{3}\circ f_{\tau}$. The action of $T_{3}$ on $\mathbb{P}^{2}$ multiplies $X_{\rho}$ by the constant $\rho^{2}$, and, in consequence, preserves the affine structure along the leaves of $\mathcal{L}_{\rho}$. This endows the leaves of $\mathcal{F}_{3}$ on the regular part of $U/T_{3}$ with an affine structure each, which moreover varies holomorphically (a foliated affine structure). The leaves of $\mathcal{L}_{\rho}$ that correspond to points in $E_{\rho}$ with trivial stabilizer under the action of $T^{\flat}_{3}$ map injectively to $\mathbb{P}^{2}/T_{3}$ as leaves of $\mathcal{F}_{3}$; their images are transcendental, have saturated neighborhoods that are injective images of covering tubes, and are without holonomy. A leaf of $\mathcal{L}_{\rho}$ that corresponds to a point in $E_{\rho}$ with non-trivial stabilizer under the action of $T_{3}$ (that is fixed by it) maps in a three-to-one ramified way to a leaf of $\mathcal{F}_{3}$ having a holonomy of order three (there are three such leaves); the affine structure in each one of these leaves is inherited from the one on $\mathbb{C}$ under the quotient by multiplication by $\rho^{2}$, with the point $0\in\mathbb{C}$ corresponding to a singular point of $\mathbb{P}^{2}/T_{3}$. It will be convenient to consider the quotient $E_{\rho}/T^{\flat}_{3}$ as an orbifold, modeled on $\mathbb{P}^{1}$, with three conic points of angle $2\pi/3$. The three leaves with holonomy correspond to the conical points of the orbifold structure for $E_{\rho}/T^{\flat}_{3}$.
@@
Let $R:\mathbb{P}^{2}/T_{3}\dashrightarrow\mathbb{P}^{2}/T_{3}$ be a birational automorphism of $\mathcal{F}_{3}$. It permutes holomorphically the Zariski-dense leaves of $\mathcal{F}_{3}$, in a holonomy-preserving way. In particular, $R$ induces a biholomorphism of $E_{\rho}/T^{\flat}_{3}$ compatible with its orbifold structure. Through its action on the three conical points, the orbifold group of biholomorphisms of $E_{\rho}/T^{\flat}_{3}$ identifies to a subgroup of the group $S_{3}$ of permutations on three symbols. This group is actually all of $S_{3}$, since it contains the maps induced by the transformations $Q$ and $S$ of Eqs. (17) and (18). In order to prove the first item of Theorem 5, we will establish that $R$ belongs to the group generated by the maps induced by $Q$ and $S$. Up to composing with an element of this group, we may suppose that $R$ acts trivially on the leaf space of $\mathcal{F}_{3}$. Our aim to establish that $R$ is the identity.
Let us show that there is a holomorphic map $\widetilde{R}:U\to U$, preserving $\mathcal{L}_{\rho}|_{U}$, such that, in restriction to $U$, $R\circ\Pi_{3}=\Pi_{3}\circ\widetilde{R}$. Let $U^{*}\subset U$ be the set formed by the leaves without holonomy, those that are not fixed by $T_{3}$. As in Section 7.1, the covering tubes around the holonomy-free leaves of $\mathcal{F}_{3}$ are mapped by $R$ to covering tubes, affinely in restriction to each leaf, as in formula (40), and, in particular, the restriction of $R$ to $U^{*}/T_{3}$ is a biholomorphism (a fact that implies that the restriction of $R$ to $U/T_{3}$ is a biholomorphism as well). Let $E_{\rho}^{*}=f_{\rho}(U^{*})$ be the complement in $E_{\rho}$ of the points fixed by $T_{3}$. The map $f_{\rho}:U^{*}/T_{3}\to E^{*}_{\rho}/T_{3}$ is a fibration with contractible leaves, and the map induced at the level of the fundamental groups is an isomorphism. The map $\Pi_{3}|_{U^{*}}:U^{*}\to U^{*}/T_{3}$ is a covering one. By the well-known criterion for the existence of lifts to covers (see, for instance, [Hat02, Prop. 1.33]), there exists a lift of $(R\circ\Pi_{3})|_{U^{*}}$, a map $\widetilde{R}^{*}:U^{*}\to U^{*}$ such that $R\circ\Pi_{3}=\Pi_{3}\circ\widetilde{R}^{*}$. We may suppose, up to composing with a suitable power of $T_{3}$, that $\widetilde{R}^{*}$ preserves each leaf of $\mathcal{L}_{\rho}$ in $U^{*}$. Let $L\subset U$ be one of the three leaves of $\mathcal{L}_{\rho}$ that has a non-trivial stabilizer under the action of $T_{3}$. Let $p\in L$ be a point that is not fixed by $T_{3}$. Consider a sufficiently small ball $B\subset U$ around $p$. In restriction to $B$, $R\circ\Pi_{3}$ is a biholomorphism onto its image, and there is thus a lift $\widetilde{r}:B\to U$ such that, in restriction to $B$, $R\circ\Pi_{3}=\Pi_{3}\circ\widetilde{r}$. Up to the action of $T_{3}$, we may suppose that $\widetilde{r}$ agrees with $\widetilde{R}^{*}$ in the intersection of the domains where each one of them is defined. In this way, $\widetilde{R}^{*}$ extends holomorphically to a neighborhood of $p$. Finally, by Hartogs’s Lemma, $\widetilde{R}^{*}$ extends to the fixed points of $T_{3}$ as well, producing the sought map $\widetilde{R}$.
@@
By Proposition 24 and its proof, $\widetilde{R}$ belongs to the flow of $X_{\rho}$. In order for such automorphism of $\mathbb{P}^{2}$ to be a lift from one of $\mathbb{P}^{2}/T_{3}$, it must normalize the group generated by $T_{3}$, this is, either $\widetilde{R}T_{3}\widetilde{R}^{-1}=T_{3}$, or $\widetilde{R}T_{3}\widetilde{R}^{-1}=T_{3}^{2}$. However, the second possibility may be discarded, for the actions of both sides of the equality on the singularities of $\mathcal{L}_{\rho}$ do not agree. Thus, $\widetilde{R}$ and $T_{3}$ commute. If $\widetilde{R}$ is given by the flow of $X_{\tau}$ in time $b$, then, in restriction to $L$, and with respect to the affine coordinate induced by $X_{\rho}$, the translation by $b$ commutes with the multiplication by $\rho^{2}$ induced by the action of $T_{3}$, but this is only possible if $b=0$. We conclude that $\widetilde{R}$ is the identity, and that $R$ is the identity as well, establishing the result.
@@ 7.3. The birational automorphisms of $\mathcal{F}_{4}$
We will now prove, along the same lines, the second point of Theorem 5, that the only non-trivial birational automorphism of $\mathcal{F}_{4}$ is the one induced by the transformation $J$ of Eq. (21).
The quotient model for $\mathcal{F}_{4}$ of Section 3.2 is given by the quotient the foliation $\mathcal{E}_{4}$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ given by (19) under the action of the order-four automorphism $T_{4}$ of Eq. (20). The foliation $\mathcal{L}_{\mathrm{i}}$ of Section 7.1 gives a birational model for $\mathcal{E}_{4}$: starting from $\mathcal{L}_{i}$ on $\mathbb{P}^{2}$, blow up the points $(1:0:0)$ and $(0:1:0)$, and then blow down the strict transform of the line originally joining them. The resulting space is $\mathbb{P}^{2}$; the foliation, $\mathcal{E}_{4}$. The automorphism $T_{4}$ of Eq. (20) corresponds to the monomial transformation associated to $T_{4}^{\prime}=\left(\begin{array}[]{rr}0&1\\ -1&0\end{array}\right)$, that, through its action on $\mathbb{H}$, fixes $\mathrm{i}$. The involution $J$ of Eq. (21) corresponds to the standard quadratic Cremona map, the monomial birational involution associated to $J^{\prime}=\left(\begin{array}[]{rr}-1&0\\ 0&-1\end{array}\right)$.
The action of $T_{4}^{\prime}$ on $E_{\mathrm{i}}$ is the order-four automorphism induced by multiplication by $\mathrm{i}$. The quotient is an orbifold modeled on $\mathbb{P}^{1}$, with three conical points: one with angle $\pi$, and two with angle $\pi/2$. Its orbifold group of biholomorphisms is generated by an involution that fixes the first point and interchanges the other two. Through its action on the leaf space, $J^{\prime}$ induces this involution.
Let $R:\mathbb{P}^{2}/T_{4}^{\prime}\dashrightarrow\mathbb{P}^{2}/T_{4}^{\prime}$ be a birational automorphism of $\mathcal{F}_{4}$. Up to composing with $J^{\prime}$, we may suppose that $R$ acts trivially on the leaf space of $\mathcal{F}_{4}$. As before, we may lift $R$ to a birational automorphism $\widetilde{R}$ of $\mathbb{P}^{2}$ that is actually holomorphic, that belongs to the flow of $X_{\mathrm{i}}$, and that commutes with $T_{4}^{\prime}$. As before, $\widetilde{R}$, and thus $R$, must be the identity. This establishes the second item of Theorem 5.
@@ 7.4. On the absence of birational automorphisms of $\mathcal{F}_{6}$
Let us now establish, in essentially the same way, the third item of Theorem 5. Consider the quotient model for $\mathcal{F}_{6}$ described in Section 3.3, and keep the objects introduced in Section 7.2. The foliation $\mathcal{E}_{6}$ is birationally equivalent to $\mathcal{L}_{\rho}$; its birational automorphism $T_{6}$, to the monomial birational transformation induced by $\left(\begin{array}[]{rr}0&1\\ -1&1\end{array}\right)$. The latter fixes $\rho$ through its action on $\mathbb{H}$. This corresponds to the biholomorphism of $E_{\rho}$ induced by multiplication by $\rho$. The quotient is an orbifold modeled on $\mathbb{P}^{1}$, with three conical points, of angles $1/2$, $1/3$ and $1/6$ times $2\pi$. It has a trivial group of biholomorphisms.
Let $R:\mathbb{P}^{2}/T_{6}\dashrightarrow\mathbb{P}^{2}/T_{6}$ be a birational automorphism of $\mathcal{L}_{\rho}$. It acts trivially on the leaf space of $\mathcal{F}_{6}$, and, as before, may be lifted to a birational automorphism $\widetilde{R}$ of $\mathbb{P}^{2}$ that belongs to the flow of $X_{\rho}$ and that commutes with $T_{6}$, and which can be shown to be the identity. This establishes the last item of Theorem 5, and finishes its proof.
Acknowledgments